OpenStudy (anonymous):

Felicia had a small bag of jelly beans. The bag contained 2 green, 10 pink, 5 red, 8 yellow, and 7 orange candies. She picked one of the jelly beans from the bag. What is the probability that Felicia picked a yellow jelly bean?

OpenStudy (anonymous):

8/total possibilities =8/(2+10+5+8+7) =8/32 =1/4